# Code Difference Guided Fuzzing for FPGA Logic Synthesis Compilers via Bayesian Optimization

### v1

@@ Abstract
Field Programmable Gate Arrays (FPGAs) play a crucial role in Electronic Design Automation (EDA) applications, which have been widely used in safety-critical environments, including aerospace, chip manufacturing, and medical devices. A critical step in FPGA development is logic synthesis, which enables developers to translate their software designs into hardware net lists, which facilitates the physical implementation of the chip, detailed timing and power analysis, gate-level simulation, test vector generation, and optimization and consistency checking. However, bugs or incorrect implementations in FPGA logic synthesis compilers may lead to unexpected behaviors in target wapplications, posing security risks. Therefore, it is crucial to eliminate such bugs in FPGA logic synthesis compilers. The effectiveness of existing works is still limited by its simple, blind mutation strategy. To address this challenge, we propose a guided mutation strategy based on Bayesian optimization called LSC-Fuzz to detect bugs in FPGA logic synthesis compilers. Specifically, LSC-Fuzz consists of three components: the test-program generation component, the Bayesian diversity selection component, and the equivalent check component. By performing test-program generation and Bayesian diversity selection, LSC-Fuzz generates diverse and complex HDL code, thoroughly testing the FPGA logic synthesis compilers using equivalent check to detect bugs. Through three months, LSC-Fuzz has found 16 bugs, 12 of these has been confirmed by official technical support.
Index Terms:
FPGA, Logic Synthesis Compiler, Equivalent Check
@@ I Introduction
In the realm of digital design and engineering, Hardware Design Language (HDL) logic synthesis tools are crucial, acting as a conduit from conceptual hardware models to executable digital circuits [1, 2, 3]. These tools allow engineers to convert high-level hardware functionality descriptions, articulated in HDL codes like VHDL or Verilog, into gate-level blueprints suitable for silicon chip fabrication. As shown as Figure 1, this conversion is essential for creating complex Integrated Circuits (ICs), encompassing CPUs, GPUs, and specialized ASICs.
@@
Within the context of Intel’s FPGA Industrial Solutions Playbook, logic synthesis assumes a critical role, enabling these devices to seamlessly bridge and integrate with legacy industrial systems. This not only facilitates the modernization of existing infrastructure but also guarantees that systems can be promptly updated or reconfigured as needs evolve—eliminating the necessity for extensive hardware modifications. This adaptability is paramount in industries characterized by rapid technological advancement and where system downtime incurs significant costs [4, 5, 6]. However, logic synthesis tools are susceptible to bugs, stemming from the complexity of implementing extensive optimizations to fulfill stringent power consumption and timing constraints [7]. If unexpected behavior occurs in FPGA logic synthesis tools, may cause further finance loss and safety accident in legacy industrial systems. Thus it is important to keep the accuracy and safety of FPGA logic synthesis tools.
@@
Recently, several methods have been developed to thoroughly test FPGA logic synthesis tools. Although they contribute significantly to testing FPGA logic synthesis tools, there remain some shortcomings. The optimal method for testing FPGA logic synthesis tools is LegoHDL [8], which generates a CPS model guided by a AST and uses HDLcoder to convert the CPS model to HDL code for FPGA logic synthesis testing. However, LegoHDL still has some disadvantage, such as low testing efficiency, and its random parameter configuration does not allow for the generation of more diverse test cases. Another testing method is Verismith [7], which uses an Abstract Syntax Tree (AST) generating method to create test cases and identify bugs. Due to the corpus’s shortcomings, Verismith can only generate less diverse and more repetitive codes. Despite its success in finding many bugs in FPGA logic synthesis compilers, Verismith’s effectiveness is still limited by its simple corpus.
We formulate our bug-finding process as an optimization problem whose goal is to maximize the difference between a variant and the seed program. By generating substantially diverse variants, we aim at testing more optimization strategies that may not be exercised otherwise. To address this issue, we propose a guided mutation strategy based on Bayesian optimization called LSC-Fuzz to detect bugs in FPGA logic synthesis compilers.
Specifically, LSC-Fuzz consists of three components: the test-program generation component, the Bayesian diversity selection component, and the equivalent check component. The test-program generation component, first collects coverage data in the seed CPS model. Then the insertion point is determined based on the data coverage collection information, and assertion statements are randomly inserted within the insertion point. Based on the coverage information of the seed model, LSC-Fuzz ensures that the inserted assertion statements are always true, and puts subsequent modules in the seed model into the true branch to ensure that they are still executed. Then LSC-Fuzz randomly select some modules to form a module sequence that conforms to the grammatical rules and add them to the newly created fake branch. The Bayesian diversity selection component first enriches the diversity of generating HDL code by adding specific timing relationship modules to the CPS model. The diversity of generated codes is then ensured by defining procedural distances and selecting LSC-Fuzz through Bayesian selection. By performing test-program generation and Bayesian diversity selection, LSC-Fuzz generates diverse and complex Verilog design file code. Finally, through the equivalent check component, if LSC-Fuzz detects inconsistencies in the output of the variant HDL code with the original HDL code, it may indicate the presence of bugs in the logic synthesis tools.
The experiments demonstrate that, within the same testing period, LSC-Fuzz significantly outperforms existing approaches, in the detection bugs of logic synthesis tools. Specifically, in three months, LSC-Fuzz has found 16 bugs, 12 of these has been confirmed by official technical support. In summary, the main contributions of this work are as follows:
  • We introduce LSC-Fuzz, a guided mutation strategy based on Bayesian optimization to uncover bugs in FPGA logic synthesis compilers. By performing test-program generation and Bayesian diversity selection, LSC-Fuzz generates diverse and complex HDL code, thoroughly testing the FPGA logic synthesis compilers using equivalent check to detect bugs.
@@
  • Extensive experiments are conducted to evaluate the bug-finding capability of LSC-Fuzz. LSC-Fuzz has detected 16 valid bugs, of which 11 bugs have been confirmed as new bugs by official technical support.
@@
  • We release LSC-Fuzz as a replication package for FPGA logic synthesis tool testing to foster reproducibility and facilitate further studies [9].
@@ II-B Illustrative Examples
We present a real confirmed bug that existing methods failed to detect. In this example, we demonstrate the effectiveness of our specific code generation approach. This bug involves a Vivado miscompilation issue, illustrated in Figure 3, which displays certain modules post-synthesis that trigger the error. Due to the complexity of the synthesized model diagram, which typically comprises thousands of modules, we focus solely on illustrating the root cause of the bug. The model generation process is outlined as follows:
Variable Profiling and Assertion Insertion: LSC-Fuzz profiles the variable valuations at all CPS points of the seed CPS model, followed by the insertion of an assertion statement. For instance, an If block is added where, based on the valuation of variables at specific CPS insertion points, the conditional predicate of this If block is enforced to always evaluate as true. Since the false condition of the If block is never executed, it introduces no side effects on the CPS model. This unused branch provides a new space to insert additional blocks, thus increasing the complexity of the generated code.
Subsystem Wrapping and HDL Conversion: Unlike procedural languages, both the true and false branches of the CPS model are encapsulated within a new subsystem. This subsystem, integral to the model, is then converted into HDL code by the HDL Coder. Throughout this process, all subsystems are translated into independent HDL code files and linked to the main file through references to ensure proper functional execution.
During synthesis of this example, the compiler crashed unexpectedly. Logically, the code within the else branch should not influence the circuit’s behavior, meaning that the generated code should be logically equivalent to the code prior to the insertion of the If block. However, this example triggered a compiler crash. The underlying issue was that the FPGA logic synthesis tool failed to appropriately manage the logic path containing the clock cycle reset signal during optimization. This failure caused the optimization algorithm to erroneously retain and execute the else branch code, which should have been disregarded. Consequently, under certain extreme timing conditions, the else branch was incorrectly selected, leading to an infinite loop and ultimately causing a crash.
Figure 3: Real FPGA logic synthesis compiler Bug found by LSC-Fuzz
III LSC-Fuzz Model
@@ III-A Overview
Figure 4: The framework of LSC-Fuzz
The framework of LSC-Fuzz is illustrated in Figure 4 and Algorithm 1. LSC-Fuzz consists of three components: the test-program generation component, the Bayesian diversity selection component, and the equivalent check component. The basic idea of LSC-Fuzz is to generate diverse and equivalent HDL test case to exercise FPGA logic synthesis compiler thoroughly with equivalent test. Specifically, the test-program generation component (lines 10–15) profiles variable valuations of the CPS model and inserts assertion statements (e.g., If/While blocks) that always evaluate to be true at arbitrary points between two blocks in the seed CPS model. Hence, the outputs of the blocks before the assertion statements can be directly passed to their subsequent blocks in the true branches to preserve the equivalence of the CPS variants. LSC-Fuzz treat the body of unexecuted branches as a new space and fill it with randomly generated modules to generate new CPS models (Figure 4(a)). The Bayesian diversity selection component (lines 16–20) uses HDL Coder convert CPS model to HDL code and use bayes selector to effectively select test case to ensure the diversity of HDL code (Figure 4(b)). It generates complex HDL code with complicated timing relationship for triggering deeper bugs. Finally, LSC-Fuzz utilizes equivalent test (lines 1–9) to validate all HDL testcase (Figure 4(c)). If the output of a generated HDL code is different from the HDL code generated by seed CPS model, it is deemed to trigger a FPGA synthesis compiler bug.
Input: HDL Coder $L$, Seed CPS model $C$, CPS parameters $P$, FPGA logic Synthesis tool $S$, HDL Code $H$, Mutant number MAX-ITER
@@
8    $H^{\prime}\leftarrow\textnormal{{TPC(}}\textnormal{\emph{C, P, L}}\textnormal{{)}}$;
9    $H^{\prime\prime}\leftarrow\textnormal{{BDS(}}\textnormal{\emph{H', S, H}}\textnormal{{)}}$;
@@
13       return ReportBug(S, $H^{\prime\prime}$, P);
@@ 18 $V\leftarrow\text{Profile}(C,P)$;
19 $a\leftarrow\text{GenAssertion}(C,P,V)$;
20 $C^{\prime}\leftarrow\text{FillSegment}(C,a)$;
@@ 25 $B\leftarrow\text{BayesSelector}(H^{\prime},H)$;
27
28    $H^{\prime\prime}\leftarrow\text{Optimize}(H^{\prime},S)$;
29 return Reported Bug $H^{\prime\prime}$;
Algorithm 1 LSC-Fuzz
@@ III-B Test-Program Generation Component
The Test-Program Generation component (Function TPC from lines 10 to 15 in Algorithm1) systematically reshapes the control-flow structure at the state-machine level through two coordinated mechanisms: structural metamorphosis via path cloning and complexity amplification through dead-state injection.
Structural Metamorphosis via Path Cloning
Rather than merely inserting surface-level conditional branches, LSC-Fuzz fundamentally reconstructs the finite state machine (FSM) topology of the seed program by cloning feasible execution paths and embedding them within conditionally inhibited transitions. This metamorphosis preserves functional equivalence while exponentially expanding the state space that the logic synthesis compiler must navigate.
Specifically, LSC-Fuzz begins by profiling the reachable state space of the seed FSM model, capturing the invariant properties at each state transition point. Using Simulink’s Signal Range Coverage tool, we extract the state variable valuations across all feasible paths under the complete input domain. For each identified stable state $s_{i}$ in the original FSM, we synthesize a metamorphic transition structure that duplicates the state’s forward paths while maintaining deterministic execution through carefully crafted guard conditions.
Figure 5: State-space expansion through metamorphic transitions in logic synthesis
The metamorphic transformation adheres to structural constraints that ensure each cloned path remains functionally equivalent to its original. For any state variable $x$ participating in the transformation, we require that $\tau(x)\neq\text{boolean}\wedge\sigma(x)\neq\text{discrete}\wedge\psi(x)\text{ remains invariant within }T$, where $\tau(x)$ denotes the variable’s data type, $\sigma(x)$ its sampling characteristics, and $\psi(x)$ the state invariant that must hold throughout the sampling period $T$. This constraint formulation ensures that our metamorphic transitions operate on substantive state variables rather than simple boolean flags, maintain continuous or variable-rate sampling to preserve temporal dynamics, and guarantee that state invariants remain intact across all execution paths.
By constructing these metamorphic state transitions where the original execution path is embedded within a deterministically true branch, we ensure that the functional behavior remains unchanged while the structural complexity increases dramatically. This forces the logic synthesis compiler to reason about significantly more complex control-flow graphs without introducing any actual behavioral modifications.
Complexity Amplification through Dead-State Injection
Within the unreachable branches created by our metamorphic transitions, LSC-Fuzz strategically injects synthesized dead states that further amplify the structural complexity. These dead states, while never executed during normal operation, force the synthesis compiler to consider additional optimization paths and state encoding strategies.
The dead-state injection process follows rigorous constraints to maintain model validity. For each injected module $x$ within the dead-state set $\mathcal{M}$, we ensure that its data type $\tau(x)\in{\mathcal{F},\mathcal{I},\mathcal{C}}$ (representing float, integer, and complex types respectively) aligns with the surrounding state variables, preventing type inconsistencies that could trigger compiler errors. Furthermore, we require that the sampling time $\sigma(x)\in{\mathcal{T}_{c},\mathcal{T}_{v}}$ maintains either continuous or variable characteristics, preserving the temporal coherence of the state machine even in unreachable regions.
Most critically, we enforce that no injected module creates combinational loops within the dead states. Formally, for any module $x\in\mathcal{M}$ and any preceding module $y\in\mathcal{P}(x)$ in its dependency chain, we verify that $\neg\mathcal{L}(x,y)$, where $\mathcal{L}$ detects potential feedback paths. This constraint prevents the introduction of unstable configurations that could cause synthesis failures or infinite optimization loops.
Through this systematic approach to dead-state injection, we populate the unreachable state space with syntactically valid but semantically inactive structures. As illustrated in Figure 5, this expanded state space compels the FPGA logic synthesis compiler to explore a substantially broader range of optimization strategies, from state encoding and minimization to resource allocation and timing optimization. The compiler must process these complex state machines without knowing a priori which states are reachable, thereby exercising optimization paths that would remain untested with simpler input programs.
This state-machine-level metamorphosis represents a fundamental departure from traditional mutation testing approaches that operate at the expression or statement level. By reshaping the control-flow topology while preserving functional equivalence, LSC-Fuzz creates test programs that stress the synthesis compiler’s ability to handle complex state machines—a critical requirement for modern hardware designs where FSM optimization directly impacts circuit performance and resource utilization.

### v2

# Structural Mutation Based Differential Testing for FPGA Logic Synthesis Compilers

@@ Abstract
Field Programmable Gate Arrays (FPGAs) play a crucial role in Electronic Design Automation (EDA) applications, which have been widely used in safety-critical environments, including aerospace, chip manufacturing, and medical devices. A critical stage in FPGA development is logic synthesis, which enables developers to translate their software designs into hardware net lists. Logic synthesis compilers can facilitate the physical implementation of the chip and ensure the software designs efficiently mapped onto standard FPGA primitives. However, bugs in FPGA logic synthesis compilers may lead to unexpected behaviors in target hardware net lists, which may cause potential security risks. Therefore, it is critical to detect and repair such bugs in FPGA logic synthesis compilers. Existing methods mainly focus on automatic fuzz testing for FPGA logic synthesis compilers. But they are limited by the redundancy and uniformity of generated test case which caused the low efficiency in FPGA logic synthesis compiler testing. To address this challenge, we propose a mutation-based method integrate with the Bayesian optimization to detect bugs in FPGA logic synthesis compilers. The method called SmootHDL, consists with three key components, the Test-program Generation Component (TPG), the Bayesian Diversity Selection Component (BDS), and the Equivalent Test Check Component (ETC). SmootHDL imitates from a set of test cases which are called original test cases. To diversify the logic of original test cases, the TPG use a control-flow mutation strategy to generate new test cases which is functional equivalent with the original test cases. Furthermore, for removing the redundancy of generated test cases, the BDS combine the Bayesian optimizer to construct a selector based on the control-flow diversity. Finally, the ETC compares the original test cases and generated test cases. Theoretically, functional equivalent test case should have same output. If there are any disagree, SmootHDL will report and reproduce it. We evaluated our method on the mainstream FPGA logic synthesis compilers. Experiments show that SmootHDL can generate more diverse test cases and detect more bugs compared with the latest testing methods for FPGA logic synthesis compilers. Over three months, SmootHDL found 16 bugs, 12 of which were confirmed by official technical support, and 6 have already been fixed.
@@ I Introduction
Logic Synthesis is a crucial stage in the development of Field Programmable Gate Arrays (FPGAs). The FPGA logic synthesis compilers play critical role as a conduit from conceptual designs to executable digital circuits [1, 2, 3]. These compilers allow engineers to convert high-level functionality descriptions, articulated in Hardware Design Language (HDL) codes like VHDL or Verilog, into gate-level net lists suitable for silicon chip fabrication. As shown as Figure 1, this conversion is essential for creating complex Integrated Circuits (ICs). Within the context of Intel’s FPGA Industrial Solutions Playbook, FPGA logic synthesis compilers ensure the ICs can be updated or reconfigured which avoid extensive hardware modifications [4, 5, 6].
@@
However, bugs in FPGA logic synthesis compilers may produce potential security risk . Thus it is important to keep the accuracy and safety of FPGA logic synthesis compilers. Recently, several methods have been proposed to detect bugs FPGA logic synthesis compilers. These methods mainly focused on using a fuzz test method to generate HDL test case for FPGA logic synthesis compilers. Although they can detect numerous bugs in FPGA logic synthesis compiler, there also remain some limitations. The optimal method for testing FPGA logic synthesis compiler is EvoHDL [7], which generates Simulink Models guided by a AST and uses HDL Coder  [8] to convert the Simulink Model to HDL test case for FPGA logic synthesis compiler testing. However, due to the limitation of Simulink Model library and the inherent optimization of HDL Coder , the diversity of generated test cases from EvoHDL was limited. Another testing method is Verismith [9], which uses an Abstract Syntax Tree (AST) generating method to create test cases and identify bugs. Despite its success in finding many bugs in FPGA logic synthesis compilers, Verismith’s effectiveness is still limited by its simple corpus.
Thus, there remains two challenges in FPGA logic synthesis compiler testing.
Challenge 1. Test Efficiency and Redundancy Existing fuzzing-based methods [7, 9], while effective in generating a large number of test cases, inherently suffer from a high degree of test case redundancy. The random nature of these approaches often leads to the repeated generation of semantically or structurally similar HDL code. This phenomenon, which we term stagnant test space exploration, results in a significant waste of computational resources and time. Instead of discovering new, valuable test cases that could uncover deep-seated compiler bugs, the testing process becomes trapped in local minima, repeatedly covering already-tested code paths.
Challenge 2. Limited Test Case Diversity Another major limitation of current methods is their reliance on simple corpus generation. Although this approach can produce syntactically valid code, the resulting test cases often exhibit semantic homogeneity. For example, these methods might generate countless test case of simple arithmetic logic, but they fail to create complex, bug-revealing structures like intricate state machines, deep pipelines, or nested loops. This lack of semantic diversity severely restricts the test’s ability to expose bugs that only manifest under complex, real-world design conditions.
To address these challenge, we proposed SmootHDL, a mutation based FPGA logic synthesis compilers testing method. Our key insight is that control-flow mutation combined with Bayesian optimization expands the test space and reduces redundancy, leading to more diverse and efficient compiler testing.
Specifically, SmootHDL consists of three main components: the Test-program Generation Component (TPG), the Bayesian Diversity Selection Component (BDS), and the Equivalent Test Check Component (ETC). SmootHDL initiates with a set of test case which composed by Simulink Models, we call them seed Simulink Model. And for the HDL test case generated by the seed Simulink Model we call it original test case. To generate diverse test cases in control flow, the TPG first collects all the Stateflow chart blocks in the seed Simulink Model. Then, the TPG extracts the state graph of the seed Simulink Model. Based on this state graph, the TPG performs structural mutations such as state duplication, path duplication, and transition expansion. Theoretically, these structural mutations should not alter the functionality of the seed Simulink Model but expand its state space. We call the new Simulink Model which has been gone through the structural mutations as mutant Simulink Model. Then the BDS transform the mutant Simulink Model into HDL test case by HDL Coder . But as we mentioned above, the inherent optimizations of HDL Coder may cause different models to converge to similar HDL structures after translation. Hence by defining timing logic program distances, the BDS use Bayesian acquisition function to build a probability matrix for further test case selection. Since structural metamorphosis produces numerous functionally equivalent but structurally different variants, whose impact on timing complexity and diversity is uncertain. Bayesian inference naturally models this uncertainty and updates beliefs from observations, enabling efficient selection of valuable variants. Finally, through the ETC , if SmootHDL detects inconsistencies in the output of the variant HDL code with the original HDL code, it may indicate the presence of bugs in the logic synthesis tools.
We evaluated our method on mainstream FPGA logic synthesis compilers. Experiments demonstrate that, within the same testing period, SmootHDL can find more bugs than existing methods. In three months of testing, we reported 16 unique bugs, 12 of which were confirmed by official developers, and 6 have already been fixed. Furthermore, SmootHDL has shown that it can generate more diverse test cases for FPGA logic synthesis compilers compared with existing methods. Finally, we verified the effectiveness of the TPG and BDS through ablation experiments.
In summary, the main contributions of this work are as follows:
  • We introduce SmootHDL, a mutation-based method for FPGA logic synthesis compiler testing. To the best of our knowledge, this is the first approach that combines control-flow mutation with a Bayesian selector for FPGA logic synthesis compiler testing.
@@
  • Extensive experiments are conducted to evaluate the bug-finding capability of SmootHDL. SmootHDL has detected 16 valid bugs, of which 12 bugs have been confirmed by official technical support and 6 have been fixed.
@@
  • We release SmootHDL as a replication package for FPGA logic synthesis tool testing to foster reproducibility and facilitate further studies [10].
@@ II-B Illustrative Examples
We present a real confirmed bug that existing methods failed to detect. In this example, we demonstrate the effectiveness of our structural mutation strategy. The bug involves a Vivado miscompilation issue, illustrated in List LABEL:lst:strategy1, which shows the synthesized modules that trigger the error. Since the synthesized model typically contains thousands of modules, we highlight only the structural root cause.
1localparam IDLE=2’b00, RUN=2’b01, RUN_DUP=2’b10;
2always @* begin
3 next_state = state;
4 case (state)
5 RUN: begin
6 next_state = RUN;
7+ if (g_inhib)
8+ next_state = RUN_DUP;
9+ if ((~rst_n) $\land$ reset_seen)
10+ next_state = RUN_DUP;
11 end
12+ RUN_DUP: next_state = RUN;
13 endcase
14end
Listing 1: Example of Metamorphosis strategies in Logic Operation Optimization
State Duplication with Inhibited Branch: SmootHDL duplicates a reachable state in the seed Simulink Model and inserts a complementary guard $g_{\mathit{inhib}}$ to control the cloned transition. Under default conditions $g_{\mathit{inhib}}=0$, the execution path remains identical to the original Simulink Model, ensuring functional equivalence. The cloned state introduces an unused structural branch in the control-flow graph, which increases the diversity and complexity of the generated HDL.
Subsystem Wrapping and HDL Conversion: During HDL conversion, the duplicated state and its inhibited transition are encapsulated into a new subsystem by HDL Coder. Each subsystem is then converted into an independent HDL file and linked into the main design. Although the inhibited branch should not affect functional behavior, the compiler is still required to process it during synthesis.
Unexpectedly, the synthesis tool crashed during optimization. Logically, the duplicated branch should have been eliminated, as its guard is never satisfied. However, the optimizer incorrectly retained this path, which involved the reset signal of the clock cycle. This mistake caused the synthesis engine to explore an invalid logic path. Under extreme timing conditions, the inhibited branch was incorrectly taken, leading to a loop in the control logic and ultimately triggering a crash.
This case demonstrates that structural mutations, such as state duplication with inhibited transitions, can expose deep bugs in FPGA logic synthesis tools that would not be discovered through simpler random fuzzing approaches.
III Methodology
In this section, we first present an overview of the SmootHDL framework and then explain its three major components: the Test Program Generation Component (TPG ) the Bayesian Diversity Selection Component (BDS ) and the Equivalent Test Check Component (ETC ).
@@ III-A Overview
Figure 3: The Framework of SmootHDL
Input: HDL Coder $L$, Seed Simulink Model $C$, Simulink parameters $P$, FPGA logic Synthesis tool $S$, HDL Code $H$, Mutant number MAX-ITER
@@
8    $C^{\prime}\leftarrow\textnormal{{TPG(}}\textnormal{\emph{C, P, L}}\textnormal{{)}}$;
9    $H^{\prime\prime}\leftarrow\textnormal{{BDS(}}\textnormal{\emph{C', H}}\textnormal{{)}}$;
@@
13       return ReportBug(S, $H^{\prime\prime}$);
@@ 18 $V\leftarrow\text{Profile}(C,P)$;
 // Extract state graph
19 $A\leftarrow\text{Extgraph}(C,P,V)$;
 // structural mutations
20 $C^{\prime}\leftarrow\text{strumut}(C,A)$;
21
22 return $C^{\prime}$;
23
24Function BDS($C^{\prime}$, $H$):
25
@@ 27 $B\leftarrow\text{BayesSelector}(H^{\prime},H)$;
29
30    $H^{\prime\prime}\leftarrow\text{Optimize}(H^{\prime})$;
31 return $H^{\prime\prime}$;
Algorithm 1 SmootHDL
An overview of the TRAGIC framework is illustrated in Figure 3. It takes Simulink Model as input and yields HDL test cases for bug detection. To diversify the Simulink Models and further diverse the control-flow of generated HDL test cases, TPG uses a control-flow mutation for the input CPS. Specifically, TPG first collects and constructs the state graph in the input CPS. Then TPG use structural mutation strategies to mutate the input Simulink Model based on the state graph. These structural mutation strategies will not change the original states of input Simulink Model but make the state flow more complex. Hence TPG makes the control flow more complex of HDL test cases generated by the input CPS. After structural mutation, new Simulink Model which mutate from the input Simulink Model will be sent to the BDS to translate to HDL test cases. Recognize that the HDL Coder may lead to homogenization for the generated test case. BDS uses a Bayesian selector to select the HDL test case based on the difference between the generated test case. It ensures that all selected HDL test cases are different in control flow structure. Finally, the ETC validate output of the HDL test case generated by the input Simulink Model and generated HDL test case via simulation. Theoretically, they should be consistent for the mutation and selection should not change the original function of the input Simulink Model. If there are any inconsistent SmootHDL report and reproduce it. We present the detailed algorithm in Algorithm 1, with lines 10-14 for the TPG module and lines 15-20 for the BDS module. Next, we elaborate on the designs for the three modules of SmootHDL.
@@ III-B Test-Program Generation Component
The TPG accepts a seed Simulink Model as input and produces structurally mutated variants that enlarge the state space while preserving functional behavior. The procedure consists of two stages: first, constructing the state graph of the input Simulink Model; second, applying structural mutation strategies on this graph to generate new Simulink Model variants.
State Graph Construction
In the first stage, the TPG detects all Stateflow chart blocks in the seed Simulink Model and extracts their states and transitions. The state graph is formally represented as
$G=(S,T),$ (1)
where $S=\{s_{1},s_{2},\ldots,s_{n}\}$ is the set of states and $T=\{t_{1},t_{2},\ldots,t_{m}\}$ is the set of transitions. Each transition $t_{k}\in T$ is expressed as
$t_{k}=(s_{i},s_{j},g),$ (2)
with $s_{i}$ as the source state, $s_{j}$ as the destination state, and $g$ as the transition guard. The graph $G$ captures the topology of the finite state machine and provides the basis for structural mutation.
Structural Mutation Strategies
Based on the constructed state graph, the TPG applies three structural mutation strategies.
State Duplication. For a state $s\in S$, a cloned state $s^{\prime}$ is created with identical actions such that
$\psi(s^{\prime})=\psi(s).$ (3)
The outgoing guard $g$ is decomposed into two complementary guards $g\land\neg g_{\mathit{inhib}}$ and $g\land g_{\mathit{inhib}}$, leading to $s$ and $s^{\prime}$ respectively. When $g_{\mathit{inhib}}=0$, the execution remains equivalent to the original FSM, while the state set expands to $S^{\prime}=S\cup\{s^{\prime}\}$.
Path Duplication. For a path from $s_{i}$ to $s_{j}$ denoted as $P(s_{i},s_{j})$, a parallel path $P^{\prime}(s_{i},s_{j}^{\prime})$ is created where $s_{j}^{\prime}$ is a clone of $s_{j}$ satisfying $\psi(s_{j}^{\prime})=\psi(s_{j})$. Only one of the two paths is active during execution, preserving functional equivalence while enlarging the transition set.
Transition Expansion. For a transition $t=(s_{i},s_{j},g)$, an intermediate state $s_{x}$ is inserted to form
$s_{i}\rightarrow s_{x}\rightarrow s_{j},$ (4)
with $\psi(s_{x})=\emptyset$. This guarantees that the composed transition preserves the semantics of $t$ while increasing the depth of the transition structure.
Through these mutations, the transformed Simulink Model maintains functional equivalence with the original while enlarging the state space. If the original model is represented as
$M=(S,T),$ (5)
and the mutated model as
$M^{\prime}=(S^{\prime},T^{\prime}),$ (6)
then the relation
$M^{\prime}\equiv M\quad\text{and}\quad|S^{\prime}|>|S|$ (7)
holds under default parameter settings. The enlarged state graph forces the logic synthesis compiler to process more complex control structures and exercise optimization strategies that remain untested with the seed Simulink Model.
Figure 4: Example of State Duplication
Consider a simple Simulink Model of a counter controller with three states $IDLE$, RUNRUN, and $DONE$. The original transitions are $IDLE\xrightarrow{start}RUN$, $RUN\xrightarrow{cnt==N}DONE$, and $DONE\xrightarrow{1}IDLE$, as shown in Figure 4 (left). After applying state duplication, a new state $DONE^{\prime}$ is introduced with $\psi(DONE^{\prime})=\psi(DONE)$. The transition from $RUN$ is split into two complementary guards, $cnt==N\land\neg g_{\mathit{inhib}}$ leading to $DONE$ and $cnt==N\land g_{\mathit{inhib}}$ leading to $DONE^{\prime}$. The resulting FSM has four states while preserving functional equivalence, as illustrated in Figure 4 (right). The functional equivalence is guaranteed by two conditions. First, the guards $cnt==N\land\neg g_{\mathit{inhib}}$ and $cnt==N\land g_{\mathit{inhib}}$ are mutually exclusive, ensuring that only one of the two transitions can be taken at runtime. Second, the cloned state $DONE^{\prime}$ is defined to have identical semantics as $DONE$, that is $\psi(DONE^{\prime})=\psi(DONE)$. As a result, for any input sequence, the execution trace of the mutated FSM can be mapped to an equivalent trace in the original FSM, thereby preserving determinism and observable behavior while introducing additional structural complexity.